OpenStudy (anonymous):

If you wanted to eliminate the y variable by adding the two equations in the following system, what could be your first step? 5x − 6y = 6 x − y = 12 Add the two equations together Multiply the second equation by −6 Multiply the first equation by −6 Multiply the second equation by 6

OpenStudy (anonymous):

Multiply second eq. by -6.

OpenStudy (anonymous):

letter b.

OpenStudy (anonymous):

the second one is okay,but i think the fourth is right too.

OpenStudy (anonymous):

@lbf check if it so.

OpenStudy (anonymous):

the 4th one is wrong,coz it end up to -6.so +-6=-12?they cannot be eliminate.

OpenStudy (anonymous):

WHY NOT 5x − 6y = 6 A x − y = 12 B A-6B=6-(-6)*12, THAT'S RIGHT TOO!

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

I MEAN LET A MINUS THE B MULTIPLIED BY 6